OpenStudy (waheguru):

For what values of K will the line intersect with the parabola at only one point 3x+k=y -4x^2-2x+3=y

OpenStudy (waheguru):

One is a parabola and one is a straight line how can I find one point of intersection?

OpenStudy (bahrom7893):

3x+k=-4x^2-2x+3 You want this to be true for only one x, thus your discriminant must be 0.

OpenStudy (bahrom7893):

Let's rewrite this equation and simplify it a bit: 4x^2+2x+3x+k-3=4x^2+5x+(k-3)

OpenStudy (bahrom7893):

Discriminant is: D = b^2-4ac = 25 - 4 * 4 * (k-3)

OpenStudy (bahrom7893):

It must be equal to 0, so: 25-16k+48=0 -16k = -48-25 -16k = -73 k = 73/16... and then just simplify and make sure i didn't make any mistakes

OpenStudy (waheguru):

the answer was 5?

OpenStudy (waheguru):

-5

OpenStudy (waheguru):

u are right its just that instead of -4x^2 its -2x^2

OpenStudy (waheguru):

2x^2 <--------------

OpenStudy (waheguru):

@bahrom7893 dude you rock the steps you showed me gave me the right answer it was my error placing a wrong digit on one of the equations ;P

OpenStudy (agent0smith):

Another way to do it, would be to find the minimum value of the parabola - the vertex. Then plug in that point into the linear equation, to find k.

OpenStudy (bahrom7893):

Cheers :D

OpenStudy (agent0smith):

Actually nm what i said... that won't necessarily work.